Startup Night 6 - 20140904 UT

Observers: JHill, BRothberg, DMiller (LBTO)
Telescope Operator: GBechetti (LBTO)
AO Support: DMiller (LBTO)
Telescope Support: JLittle (LBTO)
Software Support: MDeLaPena
Instrument Support: ESolheid (LBTO)
Engineering Support: DAshby


We were open only briefly at the beginning of the night. The remainder of the time was lost to clouds. There were a few sprinkles near 5 UT, but not any significant rain.

Doug got to try a few iterations on his new synthetic active optics reconstructor on MODS1.

LBC-Red is complaining about shuttter hardware errors again, although the shutter still seemed to be working at the end of the night.

Barry and Olga ran light leak tests on MODS1.




UT 01:38 - Waking up MODS

UT 01:42 - Sieve snap on MODS -looks OK


UT 02:10 - Enclosure open

UT 02:16 - Monitoring approaching clouds

UT 02:25 - In Observing Mode

Testing Active Optics

UT 02:30 - Doug testing new synthetic reconstructor with MODS1 in beam on SX. LBC-Red is in beam on DX. BS9166

UT 02:45 DOFPIA, /REDONLY -- struggled under bright sky with moonlit clouds

UT 02:55 monitor.ob -- LBC-Red was guiding in pseudomono parallel mode. Guiding had a significant drift in elevation. Is this caused by pointing model mismatch between the two sides? Christian pointed out later that these images show wild twisted trailing patterns -- which seems to indicate the guider fighting a rotational error.

UT 03:02 - Turning off active optics for testing (stopWFSing)


UT 03:13 - Closing Enclosure due to onslaught of clouds and poor conditions

John ran various shutter test OBs during the night.

MODS light leak tests

UT 06:06 - Running to test light leakage

UT 06:40 - Turning lower chamber flood lights on

UT 06:44 - Running to test light leakage

UT 07:20 - Running

UT 08:01 - Running -upper and lower chamber lights are on

UT 08:23 - Blue MODS stuck on 100% readout and expstatus=writing, but nothing is written to newdata. MODS update and blue fitsflush do nothing.

UT 08:24 - hit CTRL-C, aborted script - blue fitsflush did not recover image

UT 08:25 - created script which picks up at 2nd blue grating in script

UT 08:26 - MODS frozen at STATUS: Partner LBTO, Commissioing - blue side does not show nextfile

UT 08:38 - Script gets stuck at first line of Archive on script, regardless of what line that is (e.g. PARTNER, ID, PI_NAME)

UT 08:41 - Tried another script (sievesnap) and same problem

UT 08:42 - manually taking image on blue side - hanging on Starting Blue CCD Exposure

UT 08:42 - manual exposure on red side is fine (mods1r.20140904.0019)

UT 08:48 - Blue IC restarted - issued blue reset on MODS1

UT 08:50 - Red side stuck on readout 100%

UT 08:51 - Blue side test side exposure worked (mods1b.20140904.019)

UT 08:52 - Red side IC computer restarted

UT 08:54 - Red reset command issued taking test exposure on the red side (mods1r.20140904.0019)

UT 08:56 - Restarting with with Blue Grating (Dual and Red completed)

UT 09:04 - Restarting blue IC computer again

UT 09:09 - testing with script (0020 both)

UT 09:15 - edited script to reduce name size in header

UT 09:25 - smaller object names seem to have worked

UT 09:42 - starting (0029 blue, 0025 red)

UT 09:50 - stopped script - after dual grating - script had shift back to slitmask=imaging - fixed this and restarted script

UT 10:21 - lights on scripts completed

UT 10:22 - starting (0041 blue, 0033 red)

UT 10:35 - same error as - script had shift back to slitmask=imaging - aborting after 2nd exposure in dual prism

UT 10:38 - restarting with fixed script (41-43 blue are DarkMask, 33-35 red Darkmask, 44-45 blue are imaging mask, 36-37 imaging mask)

UT 11:03 - starting (55 blue, 47 red)

UT 11:16 - blue is in prism but red is still grating (blue 58,59,60 - red 50,51,52) not sure why this is happening - script specified dual prism - but red did not switch to prism mode - 61blue and 53red then switched to imaging mode

UT 11:28 - started

UT 12:03 - completed script

UT 12:04 - Putting MODS to sleep.

-- %USERSIG{BarryRothberg - 2014-09-04}%
Topic revision: r6 - 04 Sep 2014, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback